Given a standard normal distribution with mean =78 and standard deviation = 28, find P(85 < X < 102)


1
Expert's answer
2022-01-25T14:47:21-0500

Given: "\mu=78,\sigma=28"

"P(85 < X < 102)=P(X<102)-P(X<85)"

Now,

"\begin{aligned}
& P[x <102] \\
=& P\left[z <\frac{x-\mu}{\sigma}\right] \\
=& P\left[z <\frac{102-78}{28}\right] \\
=& P[z <0.857] \\
=& 0.80428 \\
\end{aligned}"

and

"\begin{aligned}
& P[x <85] \\
=& P\left[z <\frac{x-\mu}{\sigma}\right] \\
=& P\left[z <\frac{85-78}{28}\right] \\
=& P[z <0.25] \\
=& 0.59871 \\
\end{aligned}"

So,

"P(85 < X < 102)=P(X<102)-P(X<85)=0.80428-0.59871=0.20557"